sgrt(x+5)+sqrt(x-3)=4
I've tried and tried this, and just don't know how to get x. The answer is 4, but I don't know how to get there

Answer by lwsshak3(11628)

sgrt(x+5)+sqrt(x-3)=4
√(x+5)+√(x-3)=4
isolate one of the radicals
√(x+5)=4-√(x-3
square both sides
x+5=16-8√(x-3)+x-3
isolate remaining radical
8√(x-3)=8
square both sides again
64(x-3)=64
64x-192=64
64x=256
x=256/64=4